OpenStudy (deadshot):

Evaluate logv3 1 over 81?

OpenStudy (deadshot):

\[\log_{3} \frac{ 1 }{ 81 }\]

OpenStudy (mathstudent55):

81 = 3^4 1/81 = 3^(-4) log(base_b) of b^n = n